The table below shows the number of people viewing a certain movie in a movie theater. Use this table to answer questions 5-9.

Age 0-12

Age 13-17

Age 18+

Total

Male

16

4

14

34

Female

47

18

27

92

Total

63

22

41

126

What is the probability that a male person in the theatre is 13-17 years old?

Give your answer as a decimal rounded to the nearest thousandth.
Answer #1

solution

male person in the theater  13 -17 = 4 and total is 34

the probability = 4/34 =0.118

answer = 0.118